At the end of 2020, Marigold Corp. has accounts receivable of $2.65 million and an allowance for doubtful accounts of $115,50Prepare the journal entry for Marigold Corp. to write off the Sunland receivable and any journal entry necessary to reflect t(b) What is the estimated net realizable value of Marigolds accounts receivable before and after the entries above? What is If you could please answer part (b)
Answer #1
b
Before adjustments After adjustments
Net Realizable value 2534500 2534500
Book value 2650000 2634500
Workings:
Before adjustments:
Net Realizable value 2534500 =2650000-115500
After adjustments:
Net Realizable value 2534500 =(2650000-21300+5800)-(115500-21300+5800)
Book value 2634500 =2650000-21300+5800